# Rationale Behind Essay Scores: Enhancing S-LLM’s Multi-Trait Essay Scoring with Rationale Generated by LLMs

### v1

@@ Abstract
Existing automated essay scoring (AES) has solely relied on essay text without using explanatory rationales for the scores, thereby forgoing an opportunity to capture the specific aspects evaluated by rubric indicators in a fine-grained manner. This paper introduces Rationale-based Multiple Trait Scoring (RMTS), a novel approach for multi-trait essay scoring that integrates prompt-engineering-based large language models (LLMs) with a fine-tuning-based essay scoring model using a smaller large language model (S-LLM). RMTS uses an LLM-based trait-wise rationale generation system where a separate LLM agent generates trait-specific rationales based on rubric guidelines, which the scoring model uses to accurately predict multi-trait scores. Extensive experiments on benchmark datasets, including ASAP, ASAP++, and Feedback Prize, show that RMTS significantly outperforms state-of-the-art models and vanilla S-LLMs in trait-specific scoring. By assisting quantitative assessment with fine-grained qualitative rationales, RMTS enhances the trait-wise reliability, providing partial explanations about essays.
@@ 2.1 Traditional and transformer-based automated essay scoring
Traditional automated essay scoring (AES) focused on holistic scoring, predicting an overall score using handcrafted features and linear regression models Taghipour and Ng (2016); Dong and Zhang (2016); Dong et al. (2017); Cozma et al. (2018). Particularly, transformer-based models like BERT Devlin et al. (2018) significantly improved AES by capturing detailed language information Yang et al. (2020); Wang et al. (2022); Mayfield and Black (2020). These models enhanced scoring accuracy but were primarily used for holistic scoring. Extending them to multi-trait scoring is inefficient due to the need for multiple models for different traits, increasing computational costs Kumar et al. (2021); Do et al. (2024).
@@ 4.3 Baselines
To compare performance across the two datasets described earlier, we used five widely adopted vanilla S-LLMs, all encoder-decoder models Raffel et al. (2020); Chung et al. (2024); Lewis et al. (2019); Zhang et al. (2019); Beltagy et al. (2020b) designed for text generation tasks. We also included baseline models with a string kernel based model and RNN-based architectures from the referenced papers: HISK Cozma et al. (2018), STL-LSTM Dong et al. (2017) and MTL-BiLSTM Kumar et al. (2021), which align with our model’s characteristics (see Appendix A for details on each baseline). For a fair comparison with the traditional benchmark dataset (ASAP/ASAP++), we used the performance data of the four baseline models—HISK Cozma et al. (2018), STL-LSTM Dong et al. (2017), MTL-BiLSTM Kumar et al. (2021), and ArTS Do et al. (2024)—as reported in Do et al. (2024). ArTS is a model that employs the vanilla T5-base model for scoring multi-trait essays.
@@ 5.1.2 Faithfulness of rationales
Figures 5 and 6 compare the reliability (measured by QWK against human-labeled scores) of each model in predicting essay scores for the ASAP/ASAP++ dataset, using either the essays or the LLM-generated rationales (see Appendix D for the Feedback Prize dataset). Most models performed at over 80% of their essay-only performance in nearly entire traits when using the rationales, demonstrating that rationales make a meaningful contribution to S-LLMs’ essay evaluations. Given that these rationales are qualitative free-text outputs, this also indicates that they provide partial explanations for the models’ score predictions Wiegreffe et al. (2020); Jain et al. (2020); Li et al. (2023). However, as shown in Figure 5, Pegasus achieved about 40% of its essay-only performance with GPT-generated rationales and 50% with Llama-generated rationales in average, suggesting that it relies more on the intrinsic features of essays than on qualitative evaluation data.
@@ 5.2.1 Performance with ASAP/ASAP++
Since our target is to predict individual trait scores, we will focus on trait scoring rather than Overall scores. Owing to space constraints, the results are presented in Table 2 and 3. Table 2 shows model performance on the ASAP/ASAP++ dataset. Using GPT-generated rationales, RMTS applied to each of the five S-LLMs outperforms their respective vanilla versions across nearly all traits, except for Style in T5 and Word Choice and Voice in Flan-T5. T5 model shows incremental improvements with rationales, ranking first or second in every trait, including Overall. Additionally, RMTS with T5, BART, and LED outperforms the best traditional models—HISK, STL-LSTM, and MTL-BiLSTM—in every trait using GPT rationales. Although MTL-BiLSTM has a higher Overall score, the gap with RMTS-T5 is small, and RMTS focuses more on individual trait scoring.
@@ 5.2.2 Performance with Feedback Prize Dataset
To assess the broader applicability of using rationales to enhance S-LLMs in essay scoring, we conducted additional experiments with the Feedback Prize dataset, as shown in Table 4. Despite the small dataset size of 2.3K samples—about one-third of the ASAP/ASAP++ dataset—rationales improve the performance of the four S-LLMs (T5, BART, Pegasus, and LED) across most traits. However, integrating rationales does not improve the vanilla Flan-T5 model. We attribute this to the model’s inherent characteristics from instruction-fine-tuning, which may prevent it from effectively incorporating rationales with such a small dataset. Nevertheless, these results indicate that rationales generally enhance model performance, even in data-scarce environments.
@@ Appendix A Details of the baselines
T5111111https://huggingface.co/google-t5/t5-base Raffel et al. (2020): is a transformer-based model that frames NLP tasks as a text-to-text problem. In this study, we used the "google-t5/t5-base" model.
BART131313https://huggingface.co/facebook/bart-base Lewis et al. (2019): is a sequence-to-sequence models trained by corrupting text and learning to reconstruct the original text. In this study, we used the "facebook/bart-base" model.
Pegasus141414https://huggingface.co/google/pegasus-x-base Zhang et al. (2019): is designed specifically for abstractive summarization tasks, focusing on predicting whole sentences that have been masked. In this study, we used "google/pegasus-x-base" model.
LED151515https://huggingface.co/allenai/led-large-16384 Beltagy et al. (2020b): extends the transformer architecture to handle longer documents efficiently by using sparse attention mechanisms. In this study, we used "allenai/led-large-16384" model.
@@ Appendix B Length Statistics of Rationales
Appendix C Faithfulness of rationales from ASAP/ASAP++ dataset
Figure 9 shows the QWK performance of three S-LLMs—Flan-T5, Pegasus, and LED—on the ASAP/ASAP++ dataset using only LLM-generated rationales (see Figure 6 for T5 and BART).
Appendix D Faithfulness of rationales from Feedback Prize dataset
Figure 10 and 11 show the model performance in QWK on the Feedback Prize dataset using only LLM-generated rationales.
@@ Appendix E LLM Settings
For RMTS, we used GPT-3.5-Turbo and LLama3.1-8B-Instruct provided by OpenAI and Meta. GPT-3.5-Turbo was used in the form of API161616https://openai.com/index/openai-api/, and LLama 3.1-8B-Instruct was employed by utilizing the official code shared by Meta171717https://llama.meta.com/responsible-use-guide/. Regarding GPT, we performed the experiments with gpt-3.5-turbo-0125. When this study was conducted, the cost for processing input tokens with the model was $0.5 per 1M tokens, while generating output tokens was priced at $1.5 per 1M tokens. We consistently used identical hyperparameters: a temperature of 0, frequency and presence penalties both set to 0, and a Top-p value of 1 for the cumulative probability cutoff in nucleus sampling. Given that the temperature hyperparameter is set to 0, we conducted the experiment a single time. For prompts 3 to 6 of ASAP++, excerpts were excluded in both LLMs.

### v2

@@ Abstract
Existing automated essay scoring (AES) has solely relied on essay text without using explanatory rationales for the scores, thereby forgoing an opportunity to capture the specific aspects evaluated by rubric indicators in a fine-grained manner. This paper introduces Rationale-based Multiple Trait Scoring (RMTS), a novel approach for multi-trait essay scoring that integrates prompt-engineering-based large language models (LLMs) with a fine-tuning-based essay scoring model using a smaller large language model (S-LLM). RMTS uses an LLM-based trait-wise rationale generation system where a separate LLM agent generates trait-specific rationales based on rubric guidelines, which the scoring model uses to accurately predict multi-trait scores. Extensive experiments on benchmark datasets, including ASAP, ASAP++, and Feedback Prize, show that RMTS significantly outperforms state-of-the-art models and vanilla S-LLMs in trait-specific scoring. By assisting quantitative assessment with fine-grained qualitative rationales, RMTS enhances the trait-wise reliability, providing partial explanations about essays. The code is available at https://github.com/BBeeChu/RMTS.git.
@@
Seong Yeub Chu1††thanks: Both authors contributed equally to this research., Jong Woo Kim211footnotemark: 1, Bryan Wong1, Mun Yong Yi1††thanks: Corresponding author. 1Graduate School of Data Science, KAIST 2Department of Industrial & Systems Engineering, KAIST {chseye7, gsds4885, bryan.wong, munyi}@kaist.ac.kr
@@ 2.1 Traditional and transformer-based automated essay scoring
Traditional automated essay scoring (AES) focused on holistic scoring, predicting an overall score using handcrafted features and linear regression models Taghipour and Ng (2016); Dong and Zhang (2016); Dong et al. (2017); Cozma et al. (2018). Particularly, transformer-based models like BERT Devlin et al. (2018) significantly improved AES by capturing detailed language information Yang et al. (2020a); Wang et al. (2022); Mayfield and Black (2020). These models enhanced scoring accuracy but were primarily used for holistic scoring. Extending them to multi-trait scoring is inefficient due to the need for multiple models for different traits, increasing computational costs Kumar et al. (2021); Do et al. (2024).
@@ 4.3 Baselines
To compare performance across the two datasets described earlier, we used five widely adopted vanilla S-LLMs, all encoder-decoder models Raffel et al. (2020); Chung et al. (2024); Lewis et al. (2019); Zhang et al. (2019); Beltagy et al. (2020b) designed for text generation tasks. We also included baseline models with a string kernel based model and RNN-based architectures from the referenced papers: HISK Cozma et al. (2018), STL-LSTM Dong et al. (2017), MTL-BiLSTM Kumar et al. (2021), PMAES Chen et al. (2023), and PLAES Chen and Li (2024). These models align with our main task of multi-trait scoring (see Appendix A for details on each baseline). For a fair comparison with the traditional benchmark datasets (ASAP/ASAP++), we used the performance data of four baseline models—HISK Cozma et al. (2018), STL-LSTM Dong et al. (2017), MTL-BiLSTM Kumar et al. (2021), and ArTS Do et al. (2024)—as reported in Do et al. (2024), along with two additional baseline models, PMAES Chen et al. (2023) and PLAES Chen and Li (2024), as reported in their original papers. ArTS is a model that employs the vanilla T5-base model for scoring multi-trait essays.
@@ 5.1.2 Faithfulness of rationales
Figures 5 compares the reliability (measured by QWK against human-labeled scores) of each model in predicting essay scores for the ASAP/ASAP++ dataset, using either the essays or the LLM-generated rationales. Most models performed at over 80% of their essay-only performance in nearly entire traits when using the rationales, demonstrating that rationales make a meaningful contribution to S-LLMs’ essay evaluations. Given that these rationales are qualitative free-text outputs, this also indicates that they provide partial explanations for the models’ score predictions Wiegreffe et al. (2020); Jain et al. (2020); Li et al. (2023). However, as shown in Figure 5, Pegasus achieved about 40% of its essay-only performance with GPT-generated rationales and 50% with Llama-generated rationales in average, suggesting that it relies more on the intrinsic features of essays than on qualitative evaluation data.
@@ 5.2.1 Performance with ASAP/ASAP++
Since our target is to predict individual trait scores, we will focus on trait scoring rather than Overall scores. Owing to space constraints, the results are presented in Table 2 and 3. Table 2 shows model performance on the ASAP/ASAP++ dataset. Using GPT-generated rationales, RMTS applied to each of the five S-LLMs outperforms their respective vanilla versions across nearly all traits, except for Style in T5 and Word Choice and Voice in Flan-T5. T5 model shows incremental improvements with rationales, ranking first or second in every trait, including Overall. Additionally, RMTS with T5, BART, and LED outperforms the best traditional models—HISK, STL-LSTM, and MTL-BiLSTM—in every trait using GPT rationales. Although MTL-BiLSTM has a higher Overall score, the gap with RMTS-T5 is small, and RMTS focuses more on individual trait scoring (Additional performance comparisons with baselines specifically designed for overall score assessment are provided in Appendix D.2).
@@ 5.2.2 Performance with Feedback Prize Dataset
To assess the broader applicability of using rationales to enhance S-LLMs in essay scoring, we conducted additional experiments with the Feedback Prize dataset, as shown in Table 4. Despite the small dataset size of 2.3K samples—about one-third of the ASAP/ASAP++ dataset—rationales improve the performance of the four S-LLMs (T5, BART, Pegasus, and LED) across most traits. However, integrating rationales does not improve the vanilla Flan-T5 model. We attribute this to the model’s inherent characteristics from instruction-fine-tuning Chung et al. (2024), which may prevent it from effectively incorporating rationales with such a small dataset. Nevertheless, these results indicate that rationales generally enhance model performance, even in data-scarce environments.
@@ 6 Conclusion
Acknowledgements
This research was supported by Chunjae Education Inc. (Project Number G01250064).
@@ Appendix A Details of the baselines
R2BERT Yang et al. (2020b): enhances AES by fine-tuning pre-trained language models, combining regression and ranking objectives to improve scoring accuracy.
NPCR Xie et al. (2022)111111https://github.com/CarryCKW/AES-NPCR.git: employs pairwise contrastive regression to learn relative score differences between essays, thereby refining the scoring process.
PMAES Chen et al. (2023)121212https://github.com/gdufsnlp/PMAES.git: utilizes prompt-mapping contrastive learning to generalize scoring across various prompts, enhancing the model’s adaptability.
PLAES Chen and Li (2024): introduces a prompt-generalized and level-aware learning framework, improving cross-prompt AES performance by considering prompt variations and essay complexity levels.
T5131313https://huggingface.co/google-t5/t5-base Raffel et al. (2020): is a transformer-based model that frames NLP tasks as a text-to-text problem. In this study, we used the "google-t5/t5-base" model.
BART151515https://huggingface.co/facebook/bart-base Lewis et al. (2019): is a sequence-to-sequence models trained by corrupting text and learning to reconstruct the original text. In this study, we used the "facebook/bart-base" model.
Pegasus161616https://huggingface.co/google/pegasus-x-base Zhang et al. (2019): is designed specifically for abstractive summarization tasks, focusing on predicting whole sentences that have been masked. In this study, we used "google/pegasus-x-base" model.
LED171717https://huggingface.co/allenai/led-large-16384 Beltagy et al. (2020b): extends the transformer architecture to handle longer documents efficiently by using sparse attention mechanisms. In this study, we used "allenai/led-large-16384" model.
@@ Appendix B Length Statistics of Rationales
Appendix C Faithfulness of Rationales from ASAP/ASAP++ and Feedback Prize Datasets
This section presents the faithfulness analysis of rationales generated by LLMs on the ASAP/ASAP++ and Feedback Prize datasets. We compare model performance using Quadratic Weighted Kappa (QWK) scores when scoring essays versus LLM-generated rationales.
C.1 ASAP/ASAP++ Dataset
Figure 8 shows the QWK performance of five S-LLMs on the ASAP/ASAP++ dataset using only LLM-generated rationales.
C.2 Feedback Prize Dataset
Figures 10 and 9 illustrate the model performance in QWK on the Feedback Prize dataset using only LLM-generated rationales.
| Model | Overall | Con | PA | Lang | Nar | Org | Conv | WC | SF | Style | Voice | AVG |
| T5 RMTS(G) | 0.755 | 0.737 | 0.752 | 0.713 | 0.744 | 0.682 | 0.690 | 0.705 | 0.694 | 0.702 | 0.612 | 0.708 |
| T5 RMTS(G+C) | 0.754 | 0.733 | 0.757 | 0.703 | 0.740 | 0.680 | 0.696 | 0.691 | 0.688 | 0.697 | 0.604 | 0.704 |
| T5 RMTS(L) | 0.754 | 0.730 | 0.749 | 0.701 | 0.737 | 0.675 | 0.684 | 0.690 | 0.684 | 0.696 | 0.640 | 0.704 |
| T5 RMTS(L+C) | 0.752 | 0.722 | 0.746 | 0.694 | 0.734 | 0.672 | 0.679 | 0.682 | 0.682 | 0.692 | 0.621 | 0.698 |
| Model | Conv | Gram | Phr | Voc | Syn | Coh | Avg |
| T5 RMTS (G) | 0.568 | 0.550 | 0.543 | 0.430 | 0.543 | 0.498 | 0.522 |
| T5 RMTS (G+C) | 0.521 | 0.479 | 0.512 | 0.414 | 0.497 | 0.467 | 0.482 |
| T5 RMTS (L) | 570 | 0.557 | 0.535 | 0.443 | 0.534 | 0.490 | 0.522 |
| T5 RMTS (L+C) | 0.553 | 0.521 | 0.527 | 0.412 | 0.515 | 0.479 | 0.501 |
Appendix D Additional Experiments
D.1 Result of Compressed Rationales
As shown in Appendix B, some rationales were relatively long and could have hindered S-LLMs’ scoring performance. To address this, we designed a prompt to guide the LLM in removing redundant information, making the rationales more concise while retaining key details.
An analysis of the ASAP dataset showed that the average rationale length was significantly higher for the original rationales than for the shortened versions across all tested models. For instance, in the GPT-T5 setting, the original rationales averaged 277.55 words, with a maximum of 431 words, whereas the compressed versions averaged 133.04 words, with a maximum of 228 words. Similarly, in the Llama-T5 setting, original rationales averaged 298.62 words (max 500), compared to 143.35 words (max 238) for the shortened versions.
To assess the impact of rationale length, we compared T5’s performance using original versus shortened rationales. As shown in Table 5 and Table 6, models using original rationales generally outperformed those using compressed versions on average. This suggests that the original rationales provided richer information that improved the model’s ability to score essays reliably.
D.2 Comparison of Overall Score Prediction
While RMTS is inherently designed for multi-trait scoring, we adapted it to generate a single overall score for direct comparison with models that do not support multi-trait scoring.
Table 7 shows the QWK performance of our model compared to two baselines designed for overall score prediction: R2Bert Yang et al. (2020b) and NPCR Xie et al. (2022). T5 RMTS (GPT-3.5) achieves the highest average QWK score of 0.758, outperforming NPCR, the best-performing baseline, by a significant margin (7.7%).
Additionally, RMTS maintains strong performance across different prompts, demonstrating its generalization ability. These results confirm that RMTS excels not only in multi-trait scoring but also in overall score prediction.
| Model | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | AVG |
| R2Bert | 0.627 | 0.533 | 0.517 | 0.757 | 0.610 | 0.685 | 0.683 | 0.428 | 0.605 |
| NPCR | 0.719 | 0.586 | 0.650 | 0.751 | 0.766 | 0.753 | 0.683 | 0.543 | 0.681 |
| T5 RMTS(G) | 0.768 | 0.655 | 0.712 | 0.810 | 0.810 | 0.808 | 0.792 | 0.710 | 0.758 |
D.3 Ablation study of Feedback Prize Dataset
To assess the impact of trait-specific rationales, we conducted an additional experiment on the Feedback dataset (Figure 11). Using T5, Flan-T5, and BART, we fine-tuned each model following the same procedure described in Section 5.3. The results show a consistent performance drop for a given trait when its corresponding rationale is removed, confirming its importance in trait-specific assessment. Notably, models without a specific trait rationale still outperform baseline models without rationales, suggesting that rationales contribute beyond their assigned trait evaluation. However, in some cases, removing a trait rationale improved scoring performance on Flan-T5, indicating that rationale integration does not enhance its performance on the Feedback dataset, likely due to its small size, as discussed in Section 5.2.2. Although the extent of performance degradation varies across traits and models, the overall trend confirms that trait-specific rationales improve scoring accuracy, as their removal generally weakens performance.
@@ Appendix E LLM Settings
For RMTS, we used GPT-3.5-Turbo and LLama3.1-8B-Instruct provided by OpenAI and Meta. GPT-3.5-Turbo was used in the form of API181818https://openai.com/index/openai-api/, and LLama 3.1-8B-Instruct was employed by utilizing the official code shared by Meta191919https://llama.meta.com/responsible-use-guide/. Regarding GPT, we performed the experiments with gpt-3.5-turbo-0125. When this study was conducted, the cost for processing input tokens with the model was $0.5 per 1M tokens, while generating output tokens was priced at $1.5 per 1M tokens. We consistently used identical hyperparameters: a temperature of 0, frequency and presence penalties both set to 0, and a Top-p value of 1 for the cumulative probability cutoff in nucleus sampling. Given that the temperature hyperparameter is set to 0, we conducted the experiment a single time. For prompts 3 to 6 of ASAP++, excerpts were excluded in both LLMs.